# Revisiting In-context Learning Inference Circuit in Large Language Models

### v3

@@ Abstract
In-context Learning (ICL) is an emerging few-shot learning paradigm on Language Models (LMs) with inner mechanisms un-explored. There are already existing works describing the inner processing of ICL, while they struggle to capture all the inference phenomena in large language models. Therefore, this paper proposes a comprehensive circuit to model the inference dynamics and try to explain the observed phenomena of ICL. In detail, we divide ICL inference into 3 major operations: (1) Input Text Encode: LMs encode every input text (demonstrations and queries) into linear representation in the hidden states with sufficient information to solve ICL tasks. (2) Semantics Merge: LMs merge the encoded representations of demonstrations with their corresponding label tokens to produce joint representations of labels and demonstrations. (3) Feature Retrieval and Copy: LMs search the joint representations similar to the query representation on a task subspace, and copy the searched representations into the query. Then, language model heads capture these copied label representations to a certain extent and decode them into predicted labels. The proposed inference circuit successfully captured many phenomena observed during the ICL process, making it a comprehensive and practical explanation of the ICL inference process. Moreover, ablation analysis by disabling the proposed steps seriously damages the ICL performance, suggesting the proposed inference circuit is a dominating mechanism. Additionally, we confirm and list some bypass mechanisms that solve ICL tasks in parallel with the proposed circuit.
@@ 1 Introduction
In-Context Learning (ICL) (Radford et al., 2019; Dong et al., 2022) is an emerging few-shot learning paradigm: given the demonstrations $\{(x_{i},y_{i})\}_{i=1}^{k}$ consisting of [input text]-[label token] pairs and a query $x_{q}$, Language Models (LMs) take the sequence $[x_{1}][s_{1}][y_{1}]\dots[x_{k}][s_{k}][y_{k}][x_{q}][s_{q}]$ [^1] In this paper, we denote tokenization as []delimited-[][$\cdot$][ ⋅ ], and token concatenating as [][]delimited-[]delimited-[][$\cdot$][$\cdot$][ ⋅ ] [ ⋅ ]. (Fig. 1) as input and then predicts the label for $x_{q}$ by causal language modeling operation. Typically, the label tokens $y_{i}$ are preceded by and also predicted by forerunner tokens $s_{i}$ (e.g., the colon in “Label: ”). ICL has aroused widespread interest, but its underlying mechanism is still unclear.
There have been theoretical or empirical trials to characterize and explain the inference process of ICL (Xie et al., 2021; Dai et al., 2023; Wang et al., 2023; Han et al., 2023a; Jeon et al., 2024; Zheng et al., 2024). However, to capture all the operating dynamics and the observed interesting phenomenon of ICL in Large Language Models [^2] Large refers to scaled LMs trained by natural language data, such as Llama 3 (AI@Meta, 2024), contrast to simplified work that uses simple models trained and test on well-embedded input in toy models.(LLMs), a more comprehensive characterization is still necessary. Therefore, this paper proposes a unified inference circuit and measures various properties in LLMs for a conformation to the observed ICL phenomenon.
As shown in Fig. 1, we decompose ICL dynamics into 3 atomic operations on Transformer layers. Step 1: Input Text Encode: LMs encode each input text $x_{i}$ into linear representations in the hidden state of its corresponding forerunner token $s_{i}$. Step 2: Semantics Merge: For demonstrations, LMs merge the encoded representations of $s_{i}$ with the hidden state of its corresponding label tokens $y_{i}$. Step 3: Feature Retrieval and Copy: LMs retrieve merged label representations $y_{1:k}$ from Step 2 similar to the query representation $s_{q}$ in a task-relevant subspace and then merge them with the query representation. Finally, LM heads predict the label for $x_{q}$ using the label-attached query representation $s_{q}$. Steps 2 and 3 form a typical induction circuit, which is a key mechanism of ICL but only examined in synthetic scenarios (Elhage et al., 2021; Singh et al., 2024b; Reddy, 2024).
@@
We empirically find evidence for the existence of each proposed step in LLMs, and conduct more fine-grained measurements to gain insights into some phenomena observed in ICL scenarios, such as (1) positional bias: the prediction is more influenced by the latter demonstration (Zhao et al., 2021), (2) noise robustness: the prediction is not easy to be affected by demonstrations with noisy labels (Min et al., 2022), while larger models are less robust to noise (Wei et al., 2023), and (3) demonstration saturation: the accuracy improvements plateau when sufficient demonstrations are given (Agarwal et al., 2024; Bertsch et al., 2024), etc. (discussed in §5.3). Moreover, we find multiple bypass mechanisms for ICL with the help of residual connections, while the 3-phase dynamics remains dominant.
Our contributions can be summarized as: (1) We propose a comprehensive 3-step inference circuit to characterize the inference process of ICL, and find empirical evidence of their existence in LLMs. (2) We conduct careful measurements for each inference step and successfully capture a large number of interesting phenomena observed in ICL, which enhances the practicality of the proposed circuit. (3) Our ablation analysis suggests that the proposed circuit dominates, but some bypass mechanisms exist in parallel to perform ICL. We introduce some of these bypasses along with their empirical evidence.
@@ 2.1 Background & Related Works
In-context Learning. Discovered by Radford et al. (2019), ICL is an emerging few-shot learning paradigm with only feed-forward calculation in LMs. Given demonstrations $\{(x_{i},y_{i})\}_{i=1}^{k}$ composed of structured input-label pairs and a query $x_{q}$, typical ICL creates a prompt $[x_{1}][s_{1}][y_{1}]\dots[x_{k}][s_{k}][y_{k}][x_{q}][s_{q}]$, with some structural connectors (e.g. “Label: ”) including forerunner token $s_{i}$ (e.g. “: ”), as shown in Fig. 1. LMs receive such prompts and return the next token distribution, where the label token with the highest likelihood is chosen as the prediction. Explaining the principle of ICL is an unresolved research topic, although there have been some efforts on the relationship between ICL capacity and pre-training data (Li & Qiu, 2023; Singh et al., 2024b; a; Gu et al., 2023; Han et al., 2023b; Chan et al., 2022), the feature attribution of input prompt (Min et al., 2022; Yoo et al., 2022; Pan, 2023; Kossen et al., 2024), and reduction to simpler algorithms (Zhang et al., 2023; Dai et al., 2023; Xie et al., 2021; Han et al., 2023a). However, a comprehensive explanation of real-world LMs is needed to capture the operating dynamics of ICL.
Induction Circuit. Introduced by Elhage et al. (2021), an induction circuit is a pair of two cooperating attention heads from two transformer layers, where the “previous token head” writes information about the previous token to each token, and the “induction head” uses this information to identify a token that should follow each token. Such a function is implemented by two atomic operations: (1) copy the representation of the previous token $[\mathrm{A}]$ to the next token $[\mathrm{B}]$, and (2) retrieve and copy similar representations on $[\mathrm{A}]$ to the current token $[\mathrm{A}^{\prime}]$. Concisely, it performs inference in the form of $[\mathrm{A}][\mathrm{B}]\dots[\mathrm{A}^{\prime}]\Rightarrow[\mathrm{B}]$, which is similar to ICL-styled data. Therefore, this circuit has been widely used to explain the inference dynamics of ICL (Wang et al., 2023) and the emergence of ICL during pre-training (Olsson et al., 2022; Reddy, 2024; Singh et al., 2024b). Despite their valuable insights, their experiments rely on a synthetic setting: using simplified models and well-embedded (linearly separable) inputs, which differs from the practical ICL scenario using real-world LMs with many layers and complicated inputs. We bridge the gap between the synthetic and real-world settings, providing more detailed explanations for the inference dynamics of ICL based on real-world LLMs.
@@ 2.2 Experiment Settings
Models. We mainly conduct experiments on 4 modern LMs: Llama 3 (8B, 70B) (AI@Meta, 2024), and Falcon (7B, 40B) (Almazrouei et al., 2023). Unless specified, we report the results on Llama 3 70B, since its deep and narrow structure (80 layers, 64 heads) makes it easier to show hierarchical inference dynamics (discussed in §5.2). The results of other models can be found in Appendix H.2.
Datasets. We build ICL-formed test inputs from 6 sentence classification datasets, and unless specified, we report the average results on them: SST-2 (Socher et al., 2013), MR (Pang & Lee, 2005), Financial Phrasebank (Malo et al., 2014), SST-5 (Socher et al., 2013), TREC (Li & Roth, 2002; Hovy et al., 2001), and AGNews (Zhang et al., 2015).
Others. Unless specified, we use $k=4$ demonstrations in ICL inputs. For each dataset, we randomly sample $512$ test data and assign one order-fixed demonstration sequence for each test sample. About the prompt templates, etc., please refer to Appendix A.1.
@@ 3 Step 1, Input Text Encode: Semantics Encoding as Linear Representations in Hidden States
This section mainly confirms that LMs construct task-relevant and linearly separable semantic representations for every input text (demonstrations and queries) in the hidden states. Such linear representations are an important foundation for explaining the dynamics of ICL based on induction head, since attention-based feature retrieval, a key mechanism of induction head, can be easily done on linear representations. Current successful studies on simplified models and inputs (Chan et al., 2022; Reddy, 2024; Singh et al., 2024b) also assume the existence of such linear representations. Moreover, we confirm some interesting properties of the input text representations: (1) It is based on the capacity in the model weights and can be enhanced by demonstrations in context (Fig. 2 (Middle, Right)). (2) The similarity of representations is biased towards the encoding target’s position.
@@ 3.1 LLMs Encode Input Text on Forerunner Tokens in Hidden States
Forerunner Tokens Encode Input Text Representations. We plot the kernel alignment using 3 types of tokens in Fig. 2 (Left). The forerunner token, while often overlooked in previous work, produces the best input text-encoding, emerging in the early phase (layers 0-28) of the inference process, and keeping a high level to the end of inference. Interestingly, hidden states of label words are not satisfactory input text representations even with a high background value (the result at layer 0, refer to Appendix A.2.1 for details), which is a critical supplement to previous work which suggests the label tokens are pivots for collecting the information of demontrations (Wang et al., 2023).
Input Text Encoding is Enhanced by Demonstrations. We investigate the influence of contextual in on input text encoding by repeating the experiments with different $k$. As shown in Fig. 2 (Middle), when the demonstrations increase, feature alignment is enhanced, which is counterintuitive since longer preceding texts are more likely to confuse encoding targets. Such findings indicate that LMs (1) utilize contextual information to enhance the input text encoding and (2) correctly segment different demonstrations (detailed operation discussed in Appendix C).
Perplexed Texts are Encoded Worse. We investigate the correlation between kernel alignment and the perplexity of encoding targets with different $k$. Fig. 2 (Right, upper) shows a negative correlation for $k=0$, that is, LMs generate poorer encodings for more complex input text when no demonstrations are given, which can be identified as an In-weight Learning (IWL) property of the inner text encoding. While, when demonstrations are given in context (Fig. 2 (Right, lower)), the negative correlation disappears, which suggests that LMs effectively encode more complex samples with the help of demonstrations in context. More discussion about the relationship between classification performance and perplexity is in Appendix F.
@@ 3.2 Input Text Encoding is Linear and Task-Relevant but Position-biased
Figure 2: Input text encoding magnitudes (metricized by kernel alignment against feature encoded by an encoder-structured model) of hidden states in various layers in ICL scenario (The controlled experiments are results between current 6 datasets and TEE (Mohammad et al., 2018)). Left: Encoding magnitudes on hidden states from various types of token. Middle [^5] Experiments of Fig. 2 (Middle) on Llama 3 70B do not involve results on AGNews.: Encoding magnitudes with different $k$ on the forerunner tokens. Right: Encoding magnitudes in layer 24 of Llama 3 70B against the causal language modeling loss of the input text with (upper) $k=0$ and (lower) $k=8$.
@@
Input Text Encoding is Linear Separable and Task-relevant. We train a centroid classifier on hold-out 256 input samples (Cho et al., 2024), using the hidden states of a specific token in $[x_{t}][s_{t}]$ [^6] We skip the experiments on label tokens because of the leakage of ground-truth label information. from each layer and then predict the label $y_{t}$ (see Appendix A.3 for details). The results are shown in Fig. 3. The considerably high classification accuracy of the forerunner token suggests the high linear separabilities of the hidden states in the task-semantic-relevant subspaces since the centroid classifier is linear. In addition, a similar emerging trend in accuracy and kernel alignment confirms the reliability of the kernel alignment measurement.
Input Text Encoding is Biased towards Position. Ideally, the inner representations of similar queries should be highly similar regardless of their position in ICL inputs to support attention-based operations for classification. To verify this, for each encoding target, we extract the hidden states of forerunner tokens with various numbers of preceding demonstrations. We then calculate the cosine similarity between all possible pairs of the hidden states for the same target or different targets. As shown in Fig. 4, although the overall similarities on the same target are higher than on the different targets, they are both especially higher when their positions are close to each other. As to be discussed in §5.3, such positional similarity bias may lead to one flaw: demonstrations closer to the query have stronger impacts on ICL (Zhao et al., 2021; Lu et al., 2022; Chang & Jia, 2023; Guo et al., 2024). The principle of such bias is discussed in Appendix C.

### v4

@@ Abstract
In-context Learning (ICL) is an emerging few-shot learning paradigm on Language Models (LMs) with inner mechanisms un-explored. There are already existing works describing the inner processing of ICL, while they struggle to capture all the inference phenomena in large language models. Therefore, this paper proposes a comprehensive circuit to model the inference dynamics and try to explain the observed phenomena of ICL. In detail, we divide ICL inference into 3 major operations: (1) Input Text Encode: LMs encode every input text (in the demonstrations and queries) into linear representation in the hidden states with sufficient information to solve ICL tasks. (2) Semantics Merge: LMs merge the encoded representations of demonstrations with their corresponding label tokens to produce joint representations of labels and demonstrations. (3) Feature Retrieval and Copy: LMs search the joint representations of demonstrations similar to the query representation on a task subspace, and copy the searched representations into the query. Then, language model heads capture these copied label representations to a certain extent and decode them into predicted labels. Through careful measurements, the proposed inference circuit successfully captures and unifies many fragmented phenomena observed during the ICL process, making it a comprehensive and practical explanation of the ICL inference process. Moreover, ablation analysis by disabling the proposed steps seriously damages the ICL performance, suggesting the proposed inference circuit is a dominating mechanism. Additionally, we confirm and list some bypass mechanisms that solve ICL tasks in parallel with the proposed circuit.
@@ 1 Introduction
In-Context Learning (ICL) (Radford et al., 2019; Dong et al., 2022) is an emerging few-shot learning paradigm: given the demonstrations $\{(x_{i},y_{i})\}_{i=1}^{k}$ consisting of [input text]-[label token] pairs and a query $x_{q}$, Language Models (LMs) take the sequence $[x_{1}][s_{1}][y_{1}]\dots[x_{k}][s_{k}][y_{k}][x_{q}][s_{q}]$ [^1] In this paper, we denote tokenization as []delimited-[][$\cdot$][ ⋅ ], and token concatenating as [][]delimited-[]delimited-[][$\cdot$][$\cdot$][ ⋅ ] [ ⋅ ]. (Fig. 1) as input and then predicts the label for $x_{q}$ by causal language modeling operation. Typically, the label tokens $y_{i}$ are preceded by and also predicted on forerunner tokens $s_{i}$ (e.g., the colon in “Label: ”). ICL has aroused widespread interest, but its underlying mechanism is still unclear.
There have been theoretical or empirical trials to characterize and explain the inference process of ICL (Xie et al., 2021; Dai et al., 2023; Wang et al., 2023; Han et al., 2023a; Jeon et al., 2024; Zheng et al., 2024). However, to capture all the operating dynamics and many fragmented observed interesting phenomena of ICL in Large Language Models [^2] Large refers to scaled LMs trained by natural language data, such as Llama 3 (AI@Meta, 2024), contrast to simplified work that uses simple models trained and test on well-embedded input in toy models. (LLMs), a more comprehensive characterization is still necessary. Therefore, this paper tries to propose a unified inference circuit and measures various properties in LLMs for a conformation to the observed ICL phenomena.
In detail, as shown in Fig. 1, we decompose ICL dynamics into 3 atomic operations on Transformer layers. Step 1: Input Text Encode: LMs encode each input text $x_{i}$ into linear representations in the hidden state of its corresponding forerunner token $s_{i}$. Step 2: Semantics Merge: For demonstrations, LMs merge the encoded representations of $s_{i}$ with the hidden state of the corresponding label tokens $y_{i}$. Step 3: Feature Retrieval and Copy: LMs retrieve the similar semantics-merged label representations $y_{1:k}$ (from Step 2) to the query representation $s_{q}$ in a task-relevant subspace, then copy them into the query’s forerunner token representation. Finally, LM heads predict the label for $x_{q}$ using the label-attached query representation $s_{q}$. Steps 2 and 3 form a typical induction circuit, which is a key mechanism of ICL but only examined in synthetic scenarios (Elhage et al., 2021; Singh et al., 2024b; Reddy, 2024).
@@
We empirically find evidence for the existence of each proposed step in LLMs, and conduct more fine-grained measurements to gain insights into some phenomena observed in ICL scenarios, such as (1) positional bias: the prediction is more influenced by the latter demonstration (Zhao et al., 2021), (2) noise robustness: the prediction is not easy to be affected by demonstrations with wrong (noisy) labels (Min et al., 2022), while larger models are less robust to label noise (Wei et al., 2023), and (3) demonstration saturation: the accuracy improvements plateau when sufficient demonstrations are given (Agarwal et al., 2024; Bertsch et al., 2024), etc. (discussed in §5.3). Moreover, we find multiple bypass mechanisms for ICL conducted by residual connections, while the 3-phase dynamics remain dominant.
Our contributions can be summarized as: (1) We propose a comprehensive 3-step inference circuit to characterize the inference process of ICL, and find empirical evidence of their existence in LLMs. (2) We conduct careful measurements for each inference step and successfully capture a large number of interesting phenomena observed in ICL, which enhances the practicality of the proposed circuit. (3) Our ablation analysis suggests that the proposed circuit dominates, but some bypass mechanisms exist in parallel to perform ICL, and we introduce some of these bypasses along with their empirical existence evidence.
@@ 2.1 Background & Related Works
In-context Learning. Discovered by Radford et al. (2019), ICL is an emerging few-shot learning paradigm with only feed-forward calculation in LMs. Given demonstrations $\{(x_{i},y_{i})\}_{i=1}^{k}$ composed of input-label pairs and a query $x_{q}$, typical ICL creates an input $[x_{1}][s_{1}][y_{1}]\dots[x_{k}][s_{k}][y_{k}][x_{q}][s_{q}]$, with some structural connectors (e.g. “Label: ”) including forerunner token $s_{i}$ (e.g. “: ”), as shown in Fig. 1. LMs receive such inputs and return the next token distribution, where the label token with the highest likelihood is chosen as the prediction. Explaining the principle of ICL is an open question, although there have been some efforts on the relationship between ICL capacity and pre-training data (Li & Qiu, 2023; Singh et al., 2024b; a; Gu et al., 2023; Han et al., 2023b; Chan et al., 2022), the feature attribution of inputs (Min et al., 2022; Yoo et al., 2022; Pan, 2023; Kossen et al., 2024), and reduction to simpler processes (Zhang et al., 2023; Dai et al., 2023; Xie et al., 2021; Han et al., 2023a). However, a comprehensive and unified explanation on real-world LMs is still needed to capture the operating dynamics of ICL.
Induction Circuit. Introduced by Elhage et al. (2021), an induction circuit is a pair of two cooperating attention heads from two transformer layers, where the “previous token head” writes information about the previous token to each token, and the “induction head” uses the wroten information to identify a token that should follow each token. Concretely, such a function is implemented by two atomic operations in attention calculations: (1) copy the representation of the previous token $[\mathrm{A}]$ to the next token $[\mathrm{B}]$, and (2) retrieve and copy similar representations on $[\mathrm{B}]$ (copied from $[\mathrm{A}]$) to the current token $[\mathrm{A}^{\prime}]$. Concisely, it performs inference in the form of $[\mathrm{A}][\mathrm{B}]\dots[\mathrm{A}^{\prime}]\Rightarrow[\mathrm{B}]$, which is similar to the ICL diagram. Therefore, the induction circuit has been widely used to explain the inference dynamics of ICL (Wang et al., 2023) and the emergence of ICL during pre-training (Olsson et al., 2022; Reddy, 2024; Singh et al., 2024b). Despite their valuable insights, these experiments rely on a synthetic setting: using simplified models and well-embedded (linearly separable) inputs, which differs from the practical ICL scenario using real-world LMs with many layers and complicated inputs: we believe these synthetic works indicate the potential of Transformers or a locally optimal behavior, and can not demonstrate that their observation exists in scaled LLMs trained on wild data. So, in this paper, we try to bridge the gap between the synthetic and real-world settings with more detailed observations for the inference of ICL on real-world LLMs.
@@ 2.2 Experiment Settings
Models. We mainly conduct experiments on 4 modern LLMs: Llama 3 (8B, 70B) (AI@Meta, 2024), and Falcon (7B, 40B) (Almazrouei et al., 2023). Unless specified, we report the results on Llama 3 70B, since its deep and narrow structure (80 layers, 64 heads) makes it easier to show hierarchical inference dynamics (discussed in §5.2). The results of other models can be found in Appendix H.2.
Datasets. We build ICL-formed test inputs from 6 real-world sentence classification datasets, and unless specified, we report the average results on them: SST-2 (Socher et al., 2013), MR (Pang & Lee, 2005), Financial Phrasebank (Malo et al., 2014), SST-5 (Socher et al., 2013), TREC (Li & Roth, 2002; Hovy et al., 2001), and AGNews (Zhang et al., 2015).
Others. Unless specified, we use $k=4$ demonstrations in ICL inputs. For each dataset, we randomly sample $512$ test data points and assign one fixed demonstration sequence for each test sample to form a test input. About the prompt templates, etc., please refer to Appendix A.1.
@@ 3 Step 1, Input Text Encode: Semantics Encoding as Linear Representations in Hidden States
This section mainly confirms that LMs construct task-relevant and linearly separable semantic representations for every input text (demonstrations and queries) in the hidden states. Such linear representations are an important foundation for explaining the dynamics of ICL based on induction heads, since attention-based feature retrieval, a key mechanism of induction heads, can be easily done on linear representations. Current successful studies on simplified models and inputs (Chan et al., 2022; Reddy, 2024; Singh et al., 2024b) also assume the existence of such linear representations. Moreover, we confirm some interesting properties of the encoded input text representations: (1) It is based on the capacity in the model weights and can be enhanced by demonstrations in context. (2) The similarity of representations is biased towards the encoding target’s position.
@@ 3.1 LLMs Encode Input Text on Forerunner Tokens in Hidden States
Forerunner Tokens Encode Input Text Representations. We plot the kernel alignment on the 3 types of tokens in Fig. 2 (Left), where the forerunner token, while often overlooked in previous work, produces the best input text encoding, emerging in the early phase (layers 0-28) of the inference process, and keeping a high level to the end of inference. Interestingly, hidden states of label words are not satisfactory input text representations even with a high background value (the result at layer 0, refer to Appendix A.2.1), which is a critical supplement to previous work which suggests the label tokens are pivots for collecting the information of demontrations (Wang et al., 2023).
Input Text Encoding is Enhanced by Demonstrations. We investigate the influence of contextual information on input text encoding by repeating the experiments with different $k$. As shown in Fig. 2 (Middle), when the demonstrations increase, kernel alignment is enhanced, which is counterintuitive since longer preceding texts are more likely to confuse encoding targets. Such findings indicate that LMs (1) utilize contextual information to enhance the input text encoding and (2) correctly segment different demonstrations (detailed operation discussed in Appendix C).
Perplexed Texts are Encoded Worse. We investigate the correlation between kernel alignment and the perplexity of encoding targets with various $k$. Fig. 2 (Right, upper) shows a negative correlation for $k=0$, that is, LMs generate poorer encodings for more perplexed input text when no demonstrations are given, which can be identified as an In-weight Learning (IWL) property of the inner text encoding. While, when demonstrations are given in context (Fig. 2 (Right, lower)), the negative correlation disappears, which suggests that LMs effectively encode more complex samples with the help of demonstrations in context. More discussion about the correlation between classification performance and perplexity is in Appendix F.
@@ 3.2 Input Text Encoding is Linear and Task-Relevant but Position-biased
Figure 2: Input text encoding magnitudes (metricized by kernel alignment with feature encoded by an encoder-structured model) of hidden states in various layers in ICL scenario (The controlled experiments are results between current 6 datasets and TEE (Mohammad et al., 2018)). Left: Encoding magnitudes on hidden states from various types of token. Middle [^5] Experiments of Fig. 2 (Middle) on Llama 3 70B do not involve results on AGNews.: Encoding magnitudes with different $k$ on the forerunner tokens. Right: Encoding magnitudes in layer 24 of Llama 3 70B against the causal language modeling loss of the input text with (upper) $k=0$ and (lower) $k=8$.
@@
Input Text Encoding is Linear Separable and Task-relevant. We train a centroid classifier on hold-out 256 input samples (Cho et al., 2024), using the hidden states of a specific token in $[x_{t}][s_{t}]$ [^6] We skip the experiments on label tokens because of the leakage of ground-truth label information. from each layer and then predict the label $y_{t}$ (see Appendix A.3 for details). The results are shown in Fig. 3, where considerably high classification accuracy of the forerunner token suggests the high linear separabilities of the hidden states in the task-semantic-relevant subspaces since the centroid classifier is linear. In addition, a similar emerging trend in accuracy and kernel alignment confirms the reliability of the kernel alignment measurement.
Input Text Encoding is Biased towards Position. Ideally, the inner representations of similar queries should be highly similar regardless of their position in ICL inputs to support attention-based operations for classification. To verify this, for each encoding target, we extract the hidden states of forerunner tokens with various numbers of preceding demonstrations, and then calculate the cosine similarity between all possible pairs of the hidden states for the same target or different targets. As shown in Fig. 4, although the overall similarities on the same target are higher than on the different targets, they are both especially higher when their positions are close to each other. As to be discussed in §5.3, such positional similarity bias may lead to one flaw: Demonstrations closer to the query have stronger impacts on ICL (Zhao et al., 2021; Lu et al., 2022; Chang & Jia, 2023; Guo et al., 2024). The principle of such bias is discussed in Appendix C.